5 ∫ 3 (2 – x) dx

53 (2 – x) dx

Grade:12

2 Answers

Anusha
9 years ago
Ans:-53 (2-x)dx
=[2x-(x2 /2)]35
=[2(5-3)]-[(25/2)-(9/2)]
=[2(2)]-[(25-9)/2]
=4-8
=-4
Miryala Gopalakrishna
9 years ago
integration of 2-x from 3 to 5 is equal to
(2x-((x^2)/2)) from 3 to 5
=(2*5-((5^2)/2))-(2*3-((3^2)/2))
=(10-(25/2))-(6-(9/2))
=(-5/2)-(3/2)
=-8/2
=-4